OpenStudy (anonymous):

A fish swims 12.0 m in 5.0 s. It swims the first 4.0 m in 2.0 s, the next 3.0 m in 1.2 s, and the last 5.0 m in 1.8 s. What is the average velocity of the fish during the time interval between 0.0 s and 3.2 s? a. 0.0 m/s c. 4.5 m/s b. 2.2 m/s d. 2.8 m/s

OpenStudy (anonymous):

this is my last one :/ finally :)

OpenStudy (anonymous):

anyyone know?

OpenStudy (anonymous):

plezse hurry im tired

OpenStudy (anonymous):

eh? so total distance =4.0 + 3.0, total time =2.0+1.2 average speed = total distance/total time=7/3.2

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

whatever nvm

OpenStudy (anonymous):

i'm tired as hell

OpenStudy (anonymous):

yup you wrote: It swims the first 4.0 m in 2.0 s, the next 3.0 m in 1.2 s that means it took it 2.0s + 1.2s to swim 4.0m+3.0m

OpenStudy (anonymous):

I need the answer too!!!!!